OpenStudy (anonymous):

This is a hard 1 for ya'...Is a 2 part question: A certain mountain has an elevation of 19,649 ft. In 1915, the glacier on this peak covered 4 acres. By 2000 this glacier had melted to only 1 acre. A)Assuming that the glacier melted at a constant rate each year, PLEASE find the yearly rate! B)Use the answer from part A to write a linear equation that gives acreage A of this glacier t years past 1915. I need to know: * the yearly rate of change is _______acres/year *What is the acreage of the glacier t years after 1915? A = ______________

OpenStudy (anonymous):

1) Rate of melting = (Change of Acre) / (Change of Time) Change of Acre = Final Acre - Initial Acre = 1 - 4 = -3 Change of Time = Final Time - Initial Time = 2000 - 1915 + 1 = 86 * Take note that there are 86 years has passed * Rate of melting = -3 / 86 = ? acre per year 2) Amount of acre melted = Rate of melting + Number of year passed since 1915 Amount of acre melted = (-3/86) x (t) Amount of acre remaining = 4 - Amount of acre melted @bjhutchq

OpenStudy (anonymous):

I am so very confused...so what is the yearly rate of change _________acres/year?

OpenStudy (anonymous):

Rate of melting = -3 / 86 = ? acre per year That is the yearly rate of change

OpenStudy (anonymous):

ok, I am lost when I try to read the text on this example...1 more question PLEASE: what is the acreage of the glacier t years after 1915? A = _______

OpenStudy (anonymous):

Amount of acre remaining = 4 - Amount of acre melted I do not understand this line here...I am looking for A = ___________

OpenStudy (anonymous):

Amount of acre melted = Rate of melting + Number of year passed since 1915 Amount of acre melted = (-3/86) x (t) Amount of acre remaining = 4 - Amount of acre melted = A